2 marbles are selected at random from an urn which contain 4 white and 8 black marbles. what is the probability that both marbles will be white?
Answer by stanbon(75887)

2 marbles are selected at random from an urn which contain 4 white and 8 black marbles. what is the probability that both marbles will be white?
----
# of ways to select 2 white marbles: 4C2 = 6
---
# of ways to select 2 marbles randomly: 12C2 = 66
----
P(2 white) = 6/66 = 1/11
===============================
Cheers,
Stan H.
==========